OpenStudy (anonymous):

How do I find the radius of the circle: x^2+y^2-10x-16+73=0

OpenStudy (amistre64):

gotta complete the square to decipher it all

OpenStudy (lalaly):

complete the square

OpenStudy (anonymous):

complete the square

OpenStudy (anonymous):

i think you have to complete the square? has anyone suggested that yet?

OpenStudy (amistre64):

r^2 should be: -73 +(10/2)^2 + (16/2)^2

OpenStudy (amistre64):

more like, complete the "squares" :)

OpenStudy (amistre64):

the basis for completing a square such as: x^2 +6x +N , where N is some number that "completes the square" can be found in geometry. take for instance: x^2 + 6x \begin{array}c &&&6x/2\\ &x^2&x&x&x\\ &x&1&2&3\\ 6x/2&x&4&5&6\\ &x&7&8&9\\ \end{array} to fill out a complete square we need to add 9 more parts

OpenStudy (amistre64):

algebraly we complete a square by adding the square of half the number of "x" parts. x^2 + 4x + N ; where N=(4/2)^2 x^2 - 3x + N ; where N=(3/2)^2 x^2 + bx + N ; where N=(b/2)^2